OpenStudy (anonymous):

Find the derivative of log2(x^2/2)

OpenStudy (anonymous):

\[\log_2(x^2/2)\]

OpenStudy (anonymous):

My answer was \[1/(x^2/2)\ln2\] I don't believe it needs to be in simplest form

OpenStudy (loser66):

\[\log_{2} (x^2/2)= \log_{2}x^2-\log_{2}2=2\log_{2}x-1 \]

OpenStudy (loser66):

take derivative, (1)' =0 therefore just the first term needs to calculate

OpenStudy (loser66):

\[\frac{2}{xln2}\]

OpenStudy (anonymous):

The final answer is \[2/xln2\]

OpenStudy (loser66):

I think so.